OpenStudy (anonymous):

Find the equation of the line tangent to the graph of the given function at the point with the indicated x-coordinate. f(x) = (x^0.5 + 2)(x^2 + x); x = 1

hartnn (hartnn):

how much have u tried? where are u stuck?

hartnn (hartnn):

to get the equation of the line tangent u need slope of tangent, any ideas how to find slope of tangent when the function is given ?

OpenStudy (anonymous):

set the equation equal to 0?

hartnn (hartnn):

umm... its actually , finding the first derivative of the function at given point (x=1) so what u get when u derivate f(x) = (x^0.5 + 2)(x^2 + x) ?

OpenStudy (anonymous):

7.5

hartnn (hartnn):

what was that ? are u aware what is differentiation ?

OpenStudy (anonymous):

am i supposed to subsitute 1 for x?

hartnn (hartnn):

no, have u been taught differentiation ?

OpenStudy (anonymous):

very minimal

hartnn (hartnn):

what about product rule in differentiation ? because u have to use product rule here

OpenStudy (anonymous):

yes i do dy/dx = f'(x)g(x) + f(x) g'(x)

hartnn (hartnn):

good :) so now put f(x) = (x^0.5 + 2) and g(x)=(x^2 + x) and find dy/dx

OpenStudy (anonymous):

(2.5)(x^0.5+2) + (x^2+x)(3) ?

hartnn (hartnn):

umm, nopes when f(x) = (x^0.5 + 2) , what is f'(x) = ?

OpenStudy (anonymous):

not correct?

OpenStudy (anonymous):

.5x+2?

hartnn (hartnn):

nopes, derivative of x^n = n x^{n-1} and derivative of constant = 0 so u get f'(x) = 0.5 x^{-0.5} +0 got this ?

OpenStudy (anonymous):

i see, so it would be 0.5x^(-0.5) x (x^0.5 + 2)

hartnn (hartnn):

f'(x)g(x) + f(x) g'(x) = (0.5 x^{-0.5} )((x^2 + x) + f(x)g'(x) now find g'(x)

OpenStudy (anonymous):

sorry but i dont know how to find g'x

hartnn (hartnn):

you know that derivative of x^n = n x^{n-1} so what will be derivative of x^2 ?

OpenStudy (anonymous):

2

hartnn (hartnn):

what u get when u put n=2 in n x^{n-1}

OpenStudy (anonymous):

2x

hartnn (hartnn):

now its correct, and whats the derivative of x ?

OpenStudy (anonymous):

1

hartnn (hartnn):

correct so g'(x) = 2x+1 and entire derivative will be f'(x)g(x) + f(x) g'(x) = (0.5 x^{-0.5} )((x^2 + x) + (x^0.5 + 2)(2x+1) do u understand till here ? now just put x=1 in this , what u get dy/dx as ?

OpenStudy (anonymous):

so i plug in 1 for every x?

hartnn (hartnn):

yes ! and u get slope of line as dy/dx

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

so what u got the slope as ?

OpenStudy (anonymous):

i got 11.8

hartnn (hartnn):

(0.5 x^{-0.5} )((x^2 + x) + (x^0.5 + 2)(2x+1) = (0.5 1^{-0.5} )((1^2 + 1) + (1^0.5 + 2)(2(1)+1) = (0.5)(2) + (1+2)(2+1) = 1 + 3*3 = 10 got this ?

OpenStudy (anonymous):

oh i found my mistake, is that the final answer?

OpenStudy (anonymous):

or what is the final equation?

hartnn (hartnn):

that is the slope of line, m=10 you need ans equation, you know how to find the equation of line when slope and point are given ?

hartnn (hartnn):

for that u need y when x=1 y= f(x) = (x^0.5 + 2)(x^2 + x) put x=1 here, what u get y as ?

hartnn (hartnn):

so u get y= (1+2)(1+1) = 6 so the point is (x1,y1)= (1,6) and with m=10 use the point slope form of equation of line: y-y1 = m (x-x1) and u get the equation of line.

OpenStudy (anonymous):

y-6 = 10(x-1) = y = 10x+5

OpenStudy (anonymous):

i mean y=10x-4

hartnn (hartnn):

that is correct , y= 10x-4 is your final answer